# LANTERN++: Enhanced Relaxed Speculative Decoding with Static Tree Drafting for Visual Auto-regressive Models

### v1

@@ 6 Conclusion
We introduced LANTERN++, a speculative decoding framework that significantly improves acceleration in visual AR models by addressing token selection ambiguity. Our analysis highlighted the limitations of dynamic tree drafting under token selection ambiguity, where deterministic token selection leads to shallow draft trees and overly conservative acceptance probabilities, reducing acceleration efficiency. By decoupling draft selection from low-confidence predictions, LANTERN++ enables deeper accepted sequences and more effective acceleration while maintaining high image quality. However, tuning the multiplicative bound remains crucial for balancing acceleration and image quality. Future work could explore adaptive acceptance mechanisms that dynamically adjust relaxation based on token uncertainty, as well as hybrid drafting strategies that combine the strengths of static and dynamic tree structures to further improve efficiency.
@@ A.1 Visual Auto-Regressive Modeling
In visual AR models, image generation involves two main stages: generating image tokens through auto-regression and decoding the image tokens into actual image patches. In a text-to-image generation setting, given a tokenized text prompt $X_{1:N}$, the model generates a sequence of image tokens $X_{N+1:N+K}$ based on the following probability modeling:
$P(X_{N+1:N+K}\mid X_{1:N})=\prod_{\ell=1}^{K}P(x_{N+\ell}\mid X_{1:N+\ell-1}),$
where $K$ represents the total number of image tokens corresponding to the height and width of the image feature map. Since each token is predicted based solely on its preceding tokens, visual AR models require $K$ sequential steps to generate all $K$ image tokens.
Once the $K$ image tokens are generated, they are mapped to visual representation by referring to the codebook $\mathcal{C}$. Specifically, the codebook $\mathcal{C}=\{c_{1},\ldots,c_{L}\}$ consists of codes $c_{i}\in\mathbb{R}^{d}$, where each code is a latent feature used by the image encoder-decoder pair (such as VQVAE or VQGAN), and $d$ is the dimensionality of these features. As the text tokenization, since the image tokens represent indices in the codebook $\mathcal{C}$, each image token $x_{N+i}$ maps to $c_{x_{N+i}}$, which is then rearranged into a $h\times w\times d$ shaped tensor in raster-scan order (from top-left to bottom-right), for $h=H/f$ and $w=W/f$ when $f$ is down-sampling factor. After that, rearranged latent is fed into the image decoder $D:\mathbb{R}^{h\times w\times d}\rightarrow\mathbb{R}^{H\times W\times C}$ to construct an actual RGB image.
@@ A.2 Speculative Decoding
Given a sequence of tokens $X_{1:N}=(x_{1},\ldots,x_{N})$, the drafter model generates $\gamma$ draft tokens $\widehat{X}_{1:\gamma}$ as speculations for the next $\gamma$ tokens following $X_{1:N}$, based on the drafter distribution $p$. Each draft token $\hat{x}_{i}$ is sampled from the drafter distribution $p(\hat{x}_{i}|(X_{1:N},\widehat{X}_{1:i-1}))$ for each $i=1,\ldots,\gamma$, thus requiring $\gamma$ steps to generate $\gamma$ draft tokens.
@@
After obtaining the draft tokens, the concatenated sequence $(X_{1:N},\widehat{X}_{1:\gamma})$ is fed into the target model, which calculates the likelihood of each draft token $q(\hat{x}_{i}|(X_{1:N},\widehat{X}_{1:i-1})$ in parallel within a single forward pass. Each draft token $\hat{x}_{i}$ is accepted with a probability:
@@
If $\hat{x}_{i}$ is accepted, it is immediately set as $x_{N+i}=\hat{x}_{i}$. Otherwise, all subsequent tokens $\widehat{X}_{i+1:\gamma}$ are discarded, and $x_{N+i}$ is resampled from a distribution defined by $[q(\cdot|(X_{1:N},\hat{X}_{1:i-1}))-p(\cdot|(X_{1:N},\hat{X}_{1:i-1}))]_{+}$, where $[\cdot]_{+}$ denotes normalization over positive values only. As proven by Leviathan et al. (2023), this approach ensures that the distribution of the generated token sequence matches the distribution produced by the target model.

### v2

@@ 6 Conclusion
We introduced LANTERN++, a speculative decoding framework that significantly improves acceleration in visual AR models by addressing the limitations of dynamic tree drafting in LANTERN caused by token selection ambiguity. Our analysis highlighted the limitations of dynamic tree drafting under token selection ambiguity, where deterministic token selection leads to shallow draft trees and overly conservative acceptance probabilities, reducing acceleration efficiency. By decoupling draft selection from low-confidence predictions and introducing a multiplicative relaxation tailored to static tree drafting, LANTERN++ enables longer accepted sequences and more effective acceleration while maintaining high image quality. However, tuning the multiplicative bound remains crucial for balancing acceleration and image quality. Future work could explore adaptive acceptance mechanisms that dynamically adjust relaxation based on token uncertainty, as well as hybrid drafting strategies that combine the strengths of static and dynamic tree structures to further improve efficiency.
Acknowledgments
This work was partly supported by Institute for Information & communications Technology Promotion(IITP) (No.RS-2019-II190075, Artificial Intelligence Graduate School Program(KAIST), No.2022-0-00713, Meta-learning applicable to real-world problems, No.RS-2024-00457882, AI Research Hub Project) and National Research Foundation of Korea (NRF) (No.RS-2023- 00209060, A Study on Optimization and Network Interpretation Method for Large-Scale Machine Learning) grant funded by the Korea government (MSIT).
The authors acknowledge the use of ChatGPT (developed by OpenAI) for linguistic assistance, specifically for editing and clarifying the phrasing of sentences and paragraphs. ChatGPT was not involved in any aspect of the research content, including conceptual development, methodological design, or experimental design.
@@ Appendix A Additional Preliminaries
This section provides additional background necessary for understanding our methods. We adopt notation and exposition styles similar to those presented in Jang et al. (2025), with modifications and extensions where appropriate to suit our setting.
@@ A.1 Visual Auto-Regressive Modeling
Unlike other image generative models, visual AR models generate images by producing image tokens in an auto-regressive manner. Models used in this study, such as LlamaGen (Sun et al., 2024), Anole (Chern et al., 2024), and Lumina-mGPT (Liu et al., 2024), utilize VQVAE-based discrete tokenizers to decode image tokens into actual images. In the case of text-to-image models, given a tokenized text prompt $T_{1:N}$, image tokens $I_{1:K}$ are generated based on a probability distribution conditioned on the text prompt and the image tokens generated up to the previous step. This can be expressed as:
$P(I_{1:K}\mid T_{1:N})=\prod_{\ell=1}^{K}P(i_{\ell}\mid T_{1:N},I_{1:\ell-1}),$
where $K$ is determined by the image size, and since each image token is conditioned on the previous image tokens, typical visual AR models require $K$ sequential steps to generate $K$ tokens.
After generating $K$ image tokens, each token is transformed into a visual feature by referencing entries in the codebook $\mathcal{C}$. This codebook, defined as $\mathcal{C}={c_{1},\ldots,c_{L}}$, contains latent codes $c_{i}\in\mathbb{R}^{d}$, where each code serves as a feature vector in the encoder-decoder architecture (e.g., VQVAE or VQGAN), and $d$ represents the feature dimension. Similar to how text tokens correspond to indices, each image token $x_{N+i}$ points to a specific code $c_{x_{N+i}}$ in $\mathcal{C}$. These selected codes are then organized into a tensor of shape $h\times w\times d$ following a raster-scan order (left-to-right, top-to-bottom). Here, $h=H/f$ and $w=W/f$, where $f$ denotes the downsampling factor. The resulting tensor is passed to an image decoder $D:\mathbb{R}^{h\times w\times d}\rightarrow\mathbb{R}^{H\times W\times C}$ to reconstruct the final RGB image.
@@ A.2 Speculative Decoding
Speculative decoding begins with a drafter model that proposes a sequence of $\gamma$ tentative tokens, denoted as $\widehat{X}_{1:\gamma}$, as candidate continuations for an existing input sequence $X_{1:N}=(x_{1},\ldots,x_{N})$. These tokens are generated autoregressively, where each draft token $\hat{x}_{i}$ is sampled based on the previously observed tokens and prior drafts according to the distribution $p(\hat{x}_{i}|(X_{1:N},\widehat{X}_{1:i-1}))$.
@@
Once the speculative tokens have been produced, the full sequence $(X_{1:N},\widehat{X}_{1:\gamma})$ is evaluated by a higher-quality target model. Unlike the drafter, the target model processes all draft tokens simultaneously, computing the probabilities $q(\hat{x}_{i}|(X_{1:N},\widehat{X}_{1:i-1})$ in a single forward pass. Each token is retained with acceptance probability:
@@
Accepted tokens are appended to the output sequence. If a token is rejected, the remaining drafts are discarded, and a replacement for the rejected token is drawn from a normalized distribution based on the positive difference between the target and drafter probabilities:
$[q(\cdot|(X_{1:N},\hat{X}_{1:i-1}))-p(\cdot|(X_{1:N},\hat{X}_{1:i-1}))]_{+}$ (2)
As shown in the analysis by Leviathan et al. (2023), this two-stage procedure guarantees that the output sequence remains faithful to the target model’s true distribution.